OpenStudy (anonymous):

On the graph of the equation 3x + 2y = 18, what is the value of the y-intercept?

OpenStudy (radar):

Isolate the y term. 2y = -3x + 18 y = -3/2x +9 There it is in slope intercept form

OpenStudy (radar):

Or you can in the equation, 3x + 2y = 18, x=0, getting 2y = 18 y=9.